Ronold rode away from home in a car at the rate of 32 miles per hour. he walked back at the rate of 4 miles per hour. the round trip required 2 1/4 hours. how far did ronald ride?
Answer by jorel555(1290)

If x= the amount of miles driven/walked, then;
x/32+x/4=9/4;
x/32+8x/32=72/32
9x=72
x=8 miles each way!!!!!!
8/32=1/4 hr
8/4=2 hr
Total trip time = 2-1/4 hours!!!!!!